Home‎ > ‎Crew 135‎ > ‎

213-CommanderReport

Commander Report
02/13/2014
Ondrej Doule
 
We are wrapping up all experiments and are preparing the overall mission report. We solved the
malfunctionning issue on the  M-Audio recording hardware. We will be able to perform the EXP04
acoustic experiment tonight.
 
Highlights:
- Some of the drawings of the MDRS base expansion study are being finalized today (attached 800x600 samples).
- Our reporters were able today to complete more outside footage they needed for the documentary.
 
There is one ATV EVA planned for tomorrow regarding video shooting for 3 animations.
 
The Crew 135 is in great condition and mood and is running all activities according to our schedule.
    
-------------------------------------------------------------

CommanderCheckIn Report
02/13/2014
Ondrej Doule
Crew Physical Health:
All are well.

Eva Depart Time:
14:30
Eva Return Time:
17:30
Narrative Of Field Mission Results:
Filming for documentary movie about crew 135.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
Air heating system malfunction / heating is irregular
Rep Sched:
   19:00-21:00
Tomorrow Plan:
-  All data compilation - EXP04, we fixed the hardware and running recording for acoustic experiment.
- On short EVA ATV for 3D visualisation video shooting and animations around Hab
Inventory:

Support Req:
   See engineering report regarding electric heating system
Miscelleneous: